A train moving on a straight track with constant acceleration crosses a pole such that the front end of the train has a velocity of 20 m/s while crossing it and the rear end has a velocity of 30 m/s while the rear end crosses the pole. What will be the velocity of train when the mid-point of train is crossing tge pole?

Grade:

1 Answers

Sandeep Bhamoo
5 years ago
Let length of train =L, then by v²-u²=2as 900-400= 500= 2aL. Now for mid part s=L/2, v²- 400 = aL = 250 , v²=650 v=√650. Pleae do approve my abswer if correct